### v2

@@ 5.2 Results
We are interested in identifying the effects of assignment to a given postsecondary program on college graduation. College returns are typically thought of as tied to college graduation, motivating our focus on graduation-related outcomes (see kirkeboen2016 and altonjiarcidiaconomaurel2016 for further references). In addition, the extent to which students eventually graduate from the program to which they are assigned, rather than programs they switch to in later years, can be viewed as measure of performance of the assignment mechanism. However, for the average college program, initial enrollment and eventual graduation are far from being perfectly correlated (OECD 2019, larroucaurios2021). Therefore, we first look at whether assignment to a given program increases one’s probability of graduating from that same program. We also consider graduation from a top university as another outcome of interest. The choice of this second outcome is in line with the literature on returns to education and college choice, which highlights the role of institution quality in driving returns (see kirkeboen2016 and altonjiarcidiaconomaurel2016 for more references). In our setting, the set of “top” universities consists of Pontificia Universidad Católica de Chile in Santiago (PUC Santiago; hereafter PUC) and Universidad de Chile (UChile).
We present two types of results. First, we discuss the economic implications of our results by focusing on a single program $j$ and popular next-best programs $k$ associated with it. Then, we further illustrate the implementation of our method by estimating bounds on the treatment effects of interest for a large number of pairs $(j,k)$ and presenting statistics across these pairs. The bounds we present throughout this section represent estimates—not inference bounds. Inference is beyond the scope of this paper. [^11] Beyond the fact that the admission cutoffs are estimated, the fact that the $\delta$s used in the construction of bounds are partially identified must be taken into account when confidence bands are derived—a task that we leave for future research.
As noted earlier, the identification proposed in this paper resembles a sharp RD design and is therefore well suited for studying the effects of assignment to a program on graduation-related outcomes. In contrast, estimating the effects of graduation from a program on earnings typically involves a fuzzy design, which we study in separate ongoing work. [^12] At the time of writing of the present paper, we do not have access to earnings data and therefore cannot provide bounds here on the intention-to-treat (ITT) effects on earnings.
@@
We use medicine at PUC in Santiago as our illustrative program $j$. The choice of this program is driven by two considerations. First, to obtain reasonably precise estimates, we need sufficient observations, and this program is popular. Second, assignment to the program is interesting in itself in regard to our graduation outcomes of interest. The chosen program is a very popular program at a selective university, such that it is particularly relevant to look at its associated outcomes. Descriptive statistics about the program and its applicants are shown in the right panels of Tables 3 and 4 in Appendix C. We estimate the effects on graduation of being assigned to medicine at PUC instead of being assigned to a second-best option. We consider the five most popular second-best options to medicine at PUC—namely, medicine at UChile, medicine at University of Concepción, medicine at University of Santiago de Chile, science at PUC, and engineering at PUC. As a simple procedure to construct a sample local to the admission cutoff $c_{j}$ of interest (medicine at PUC), we use a 30-point bandwidth on either side of $c_{j}$. The resulting estimation sample is therefore invariant across our two outcomes, that is, graduation from medicine at PUC and graduation from a top university.
Figure 2: Estimated effects of assignment to medicine at PUC, under the SPO+UMAS assumptions
The figure shows estimated bounds for the effect of assignment to medicine at PUC relative to the effect of assignment to the five most frequent next-best alternatives on students’ probability of graduating from medicine at PUC (top panel) and their probability of graduating from a top university, that is, PUC or UChile (bottom panel). The bounds on this picture are estimated under the SPO and UMAS assumptions and with a 30-point bandwidth on either side of the assignment cutoff of the program of interest (medicine at PUC). Note that the bounds represent bounds on estimates, not inference bounds.
Given our choice of bandwidth, in estimating our bounds for the treatment effects of interest, the next step of the exercise is to recover local preferences $Q_{j}$ for each student. Absent strategic behavior, students’ next-best alternatives are straightforwardly given by their application lists—direct observation of $P$ and placement scores allows to compute cutoffs, budget sets, and finally $P_{j}$, which equals $Q_{j}$ under truth-telling. When we account for strategic behavior, $Q_{j}$ is not fully revealed by the observed $P_{j}$. Section 4.1 shows how the set of possible $Q_{j}$s, $\bm{\mathrm{Q}}_{j}$, can be constructed under different sets of assumptions. The weak partial order (WPO) assumption implies that any submitted list is a subset of acceptable programs ordered just as in the student’s true preferences. The strong partial order (SPO) assumption implies, in addition, that any student submitting a list strictly shorter than permitted reveals the student’s complete true list of acceptable schools. Under the SPO assumption, true local preferences $Q_{j}$ are observed for students submitting strictly fewer than eight choices (in the case of our empirical application). For students ranking eight choices, $Q_{j}$ is observed only if $\mathbf{Q}_{j}$ is a singleton. The UMAS assumption (Assumption 5) can be used in combination with WPO and SPO to reduce the size of $\mathbf{Q}_{j}$ for those students for whom it is not already a singleton (Corollary 1). We first present the results obtained under our strongest set of assumptions (SPO in combination with UMAS) in Figure 2 and then discuss how the estimated bounds vary as we impose weaker assumptions (Figure 3).
Figure 2 shows the estimated bounds on the effect on graduation of being assigned to medicine at PUC instead of being assigned to a next-best alternative to medicine at PUC. The top panel shows the estimated effects on the probability of graduating from medicine at PUC Santiago and the bottom panel on the probability of graduating from any program at a top university.
Focusing first on the top panel, we can make several comments. First, reassuringly, the estimates show that assignment to the case-study program, medicine at PUC Santiago, tends to increase students’ probability of graduating from that program. Indeed, our bounds exclude zero and identify a positive treatment effect in all five cases. Our bounds also reveal that the effect is heterogeneous across the next-preferred alternatives. This is shown by the fact that, in some cases, the bounds do not overlap across the next-preferred alternatives. For instance, our bounds suggest that admission to medicine at PUC Santiago increases the probability of graduating from that program to a larger extent for students whose next-preferred alternative is medicine at University of Santiago de Chile than for those whose next-preferred alternative is medicine at University of Concepción, engineering at PUC, or science at PUC.
In the bottom panel, our estimates show that admission to the case-study program, medicine at PUC Santiago, tends to increase students’ probability of graduating from a top university (i.e., PUC or UChile). For this outcome, our bounds exclude zero and identify a positive treatment effect in four of the five cases. Notably, though, the finding of a positive effect does not hinge on students’ next-preferred option not being a top university. In fact, we find that admission to medicine at PUC Santiago increases the probability of graduating from a top university for students whose next-preferred option is not at a top university (such as medicine at University of Santiago de Chile or University of Concepción) and for students whose next-preferred option is at PUC, too (e.g., in science or engineering). Similarly to those for the previous outcome, our estimates reveal that the effect is heterogeneous across the next-preferred alternatives, as the bounds do not overlap across a number of next-preferred options.
Figure 3: Estimated effects of assignment to medicine at PUC, across assumptions
The figure shows estimated bounds for the effect of assignment to medicine at PUC relative to the effect of assignment to the five most frequent next-best alternatives on students’ probability of graduating from medicine at PUC (top panel) and their probability of graduating from a top university, that is, PUC or UChile (bottom panel). For each next-best alternative, four sets of bounds are shown: from left to right and from darkest to lightest gray, the bounds pictured are derived under SPO in combination with UMAS, SPO alone, WPO in combination with UMAS, and WPO alone. All sets of bounds are estimated with a 30-point bandwidth on either side of the assignment cutoff of the program of interest (medicine at PUC). Note that the bounds represent bounds on estimates, not inference bounds.
We interpret these results as evidence that students’ preferences matter for their graduation outcomes, suggesting that preferences are correlated with ability or effort choices.
Figure 3 shows how the estimated bounds change across sets of assumptions for our two outcomes of interest. Just as Figure 2 does, Figure 3 shows the estimated effects of being assigned to medicine at PUC relative to the effects of being assigned to one of the five most frequent next-preferred alternatives. For each next-best alternative, four sets of bounds are shown: from left to right and from darkest to lightest gray, the bounds pictured are derived under SPO in combination with UMAS (thus coinciding with those pictured in Figure 2), SPO alone, WPO in combination with UMAS, and WPO alone. As we relax the assumptions we use to construct $\bm{\mathrm{Q}}_{j}$, the estimated bounds become wider. Under SPO alone, we are still able to draw the same conclusions as above—namely, that students’ preferences matter for their graduation outcomes, suggesting that preferences are correlated with ability or effort choices. Under SPO, we do find that assignment to medicine at PUC has a strong positive effect on the probability of graduating from the program for those students whose next-preferred alternative is any of the five we consider and that the magnitude of this effect differs across the next-preferred alternatives. For instance, the SPO bounds for the treatment effect on graduation from medicine at PUC for those whose next-preferred option is science at PUC do not overlap with the bounds for those whose next-preferred option is engineering at PUC or medicine at University of Santiago. The conclusions relative to our second outcome of interest—graduation from a top university—also continue to hold under SPO alone. The bounds derived under WPO, with and without UMAS, are much wider than those derived under SPO, demonstrating the identifying power of the SPO assumption. In particular, we do not recover heterogeneity in the treatment effects across the next-preferred alternatives. When we use WPO in combination with UMAS, our bounds still recover the positive sign of the treatment effect on each outcome for three next-preferred alternatives.
Table 1: Local samples and estimated $\underline{\delta}_{j,k}$s for $j=$ medicine at PUC
(1) (2) (3) (4)
SPO SPO WPO WPO
UMAS UMAS
Next-preferred $k$: medicine at UChile
Observations with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 772 772 964 964
Share listing $<$ 8 choices .90 .90 .92 .92
$\underline{\delta}_{j,k}^{+}$55 .32 .07 .03
$\underline{\delta}_{j,k}^{-}$53 .53 .03 .03
Next-preferred $k$: medicine at U.Concepción
Observations with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 93 99 437 627
Share listing $<$ 8 choices .59 .55 .91 .92
$\underline{\delta}_{j,k}^{+}$ 1 .46 .1 .03
$\underline{\delta}_{j,k}^{-}$ 1 1 .92 .92
Next-preferred $k$: medicine at U.Santiago
Observations with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 63 75 161 475
Share listing $<$ 8 choices .69 .58 .88 .93
$\underline{\delta}_{j,k}^{+}$73 .52 .06 .03
$\underline{\delta}_{j,k}^{-}$84 .84 .16 .16
Next-preferred $k$: science at PUC
Observations with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 81 102 224 477
Share listing $<$ 8 choices .85 .67 .94 .93
$\underline{\delta}_{j,k}^{+}$92 .92 .49 .49
$\underline{\delta}_{j,k}^{-}$ 1 1 .71 .71
Next-preferred $k$: engin. at PUC
Observations with $\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset$ 97 113 1142 1362
Share listing $<$ 8 choices .51 .44 .95 .95
$\underline{\delta}_{j,k}^{+}$ 1 .68 .22 .02
$\underline{\delta}_{j,k}^{-}$88 .88 .42 .42
@@
To illustrate how the different assumptions affect the derivation of estimates, Table 1 provides descriptive statistics on the sample used in the estimation of the bounds in each case and the implied (estimated) $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ for each pair $(j,k)$ and each set of assumptions. The equations in Proposition 5 make clear how the different assumptions affect the values taken by our bounds. The different assumptions are used to construct $\bm{\mathrm{Q}}_{j}$ for each individual and therefore determine the conditioning set $\{\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ and the values we estimate for $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$. Focusing first on the comparison of SPO and WPO, we note that an individual’s set $\bm{\mathrm{Q}}_{j}$ is larger when constructed under WPO than when constructed under SPO. This implies that the conditioning set $\{\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ is larger under WPO than it is under SPO—as seen in Table 1 when we compare the first row of each panel across Columns (1) and (3) or (2) and (4). Given a fixed bandwidth around the cutoff $c_{j}$, the set $\{\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ being larger under WPO than under SPO means that the denominator in the equations defining $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ is larger under WPO than under SPO. In addition, as the set $\bm{\mathrm{Q}}_{j}$ is larger when constructed under WPO than when constructed under SPO, the share of individuals for whom $Q_{j}$ is found to be a singleton is lower under WPO than it is under SPO. This means that the numerator in the equations defining $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ is smaller under WPO than under SPO. Overall, this leads to $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ being smaller under WPO than under SPO, as seen in Table 1. The increase in $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ as we impose SPO instead of WPO unambiguously tends to reduce the width of the bounds. The exact extent to which the width of the bounds shrinks as we impose SPO instead of WPO also depends on the joint distribution of the outcome $Y$ and $\bm{\mathrm{Q}}_{j}$, however. Turning to the role of the UMAS assumption (UMAS hereafter), we note that as we impose UMAS, certain elements are eliminated from individuals’ sets $\bm{\mathrm{Q}}_{j}$. As a consequence, following the same logic as above, $\{\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ is larger absent UMAS then under UMAS, and $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ are smaller absent UMAS than under UMAS. We can see this in Table 1 by comparing Columns (1) to (2) or (3) to (4). Again, the increased values of $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ are the key force that tends to reduce the width of the bounds as UMAS is imposed.
Results for a larger set of pairs.

### v3

@@ 5.2 Results
We are interested in identifying the effects of assignment to a given postsecondary program on college graduation. College returns are typically thought of as tied to college graduation, motivating our focus on graduation-related outcomes (see Kirkeboen et al. (2016) and Altonji et al. (2016) for further references). In addition, the extent to which students eventually graduate from the program to which they are assigned to (or from programs their assigned program is a pathway to) can be viewed as measure of performance of the assignment mechanism. However, for the average college program, initial enrollment and eventual graduation are far from being perfectly correlated (OECD 2019, Larroucau and Rios (2021)). We also consider graduation from a top university as another outcome of interest. The choice of this second outcome is in line with the literature on returns to education and college choice, which highlights the role of institution quality in driving returns (see Kirkeboen et al. (2016) and Altonji et al. (2016) for more references). In our setting, the set of “top” universities consists of Pontificia Universidad Católica de Chile in Santiago (PUC Santiago; hereafter PUC) and Universidad de Chile (UChile).
[…]
5.2.1 Average Structural Functions and the Importance of Preferences for Graduation Outcomes
Setup.
[…]
As a simple procedure to construct a sample local to each of the admission cutoffs $c_{j}$ of interest, we use a 30-point bandwidth on either side of $c_{j}$. [^14] The resulting estimation sample is therefore invariant across outcomes once the pair $(j,k)$ is fixed. Given our choice of bandwidth, the next step of the exercise is to recover local preferences $Q_{j}$ for each student at each cutoff $c_{j}$. In the absence of strategic behavior, these preferences can be directly inferred from application lists —direct observation of $P$ and placement scores allows to compute cutoffs, budget sets, and finally $P_{j}$, which equals $Q_{j}$ under truth-telling. However, under strategic behavior, $Q_{j}$ is not fully revealed by the observed $P_{j}$. Section 4.1 shows how to construct the set of possible $Q_{j}$s, that is, the set $\bm{\mathrm{Q}}_{j}$, under various assumptions. The weak partial order assumption (hereafter WPO) implies that any submitted list is a subset of acceptable programs ordered just as in the student’s true preferences. The strong partial order assumption (hereafter SPO) adds that students who submit shorter-than-maximum lists reveal their full set of acceptable programs. Under SPO, we fully observe $Q_{j}$ for students who rank strictly fewer than the maximum number of programs (eight, in the case of our empirical application). For students ranking eight choices, $Q_{j}$ is observed only if the set $\mathbf{Q}_{j}$ comes out as a singleton. The UMAS assumption (Assumption 5, hereafter UMAS) can be used in combination with WPO and SPO to refine $\mathbf{Q}_{j}$ for the non-singleton cases (Corollary 1). Other assumptions that make sense in the empirical context of interest can be used to further refine SPO and WPO bounds. In a context with a large number of alternatives and a small cap $K$, like in Chile’s college application system, the $\bm{\mathrm{Q}}_{j}$ constructed under WPO will likely be large. As discussed in Section 4.1, additional context-specific assumptions can be combined with UMAS to reduce the size of $\bm{\mathrm{Q}}_{j}$. In the applied literature, fields of study are widely recognized as key drivers of student preferences—and of the heterogeneity in preferences across students—for post-secondary programs (Altonji et al., 2016). Building on this insight, we propose the following fields assumption:
Assumption 8.
Suppose the set of all programs is partitioned into fields of study. Let $\mathcal{F}$ be a set of programs in an arbitrary field of study in this partition. Consider any cutoff $c_{j}$. Let $\omega$ be a student with score $S_{j}$ near $c_{j}$. If no program from field $\mathcal{F}$ appears in the reported list $P(\omega)$, then no program from $\mathcal{F}$ can be part of the student’s true local preference at $c_{j}$, i.e.,
$\forall\mathcal{F},P(\omega)\cap\mathcal{F}=\emptyset\Rightarrow Q_{j}(\omega)\notin\mathcal{F}\times\mathcal{F}.$
In practice, we partition programs in five fields of study: Health/Medicine; STEM; Economics/Business; Law; and Other. So for instance, if a student’s reported preferences include programs in Medicine, STEM, and Economics only, Assumption 8 excludes programs in Law and Other from their $\bm{\mathrm{Q}}_{j}$. For any program $\ell\in\mathcal{J}$, let $\mathcal{F}(\ell)$ denote the set of all programs in the field of study of program $\ell$. Researchers may apply Assumption 8 to refine the set $\bm{\mathrm{Q}}_{j}(\omega)$ of student $\omega$ as follows:
$\bm{\mathrm{Q}}_{j}^{*}(\omega)=\bm{\mathrm{Q}}_{j}(\omega)\setminus\left\{(\ell,\ell^{\prime})\in\mathcal{J}\times\mathcal{J}^{0}:P(\omega)\cap\mathcal{F}(\ell)=\emptyset~~\text{or}~~P(\omega)\cap\mathcal{F}(\ell^{\prime})=\emptyset\right\},$
where we adopt the convention that $\mathcal{F}(0)=\mathcal{J}$.
Figure 3 reports estimated bounds on the probability of graduating if assigned to Bachillerato de Ingreso Común at UChile across local first-best programs $j$, for students whose local next best is Bachillerato de Ingreso Comun at UChile. The left panel shows the probability of graduation from the local first-best program; the right panel shows the probability of graduation from a top university.
Figure 3: Graduation Rates if Assigned to Bachillerato de Ingreso Comun at UChile
The figure shows estimated bounds for $\mathbb{E}[Y(k)|Q_{j}=(j,k),S_{j}=c_{j}]$ (y-axis) with $k=$ Bachillerato de Ingreso Comun at UChile for several local first-best programs $j$ (x-axis), two binary graduation outcomes (left and right panels), and four different behavioral assumptions used to construct $\bm{\mathrm{Q}}_{j}$ (four bars for each $j$ in each panel). The outcome in the right panel is graduation from a top university (PUC or UChile) and graduation from the local first-best choice $j$ in the left panel. The four behavioral assumptions are, from left to right, SPO, WPO in combination with UMAS and fields, WPO in combination with UMAS, and WPO alone. The average parameter denotes the probability of graduating (from a top university or from program $j$) among those individuals with $S_{j}=c_{j}$ and $Q_{j}=(j,k)$ if they were all assigned to program $k$. All sets of bounds are estimated with a 30-point bandwidth on either side of the assignment cutoff $c_{j}$ of the program of interest. Note that these are estimated bounds and not inference bounds.
@@ 5.2.1 Average Structural Functions and the Importance of Preferences for Graduation Outcomes
We begin by discussing results under WPO with UMAS and the fields assumption, which we take as our preferred specification. This combination retains the minimal behavioral restrictions of WPO while incorporating two intuitive and contextually grounded assumptions—UMAS and fields—that meaningfully refine the set of plausible preferences and yield relatively tight bounds. For both outcomes, we find heterogeneous effects across different local first-best programs. This is evident from the non-overlapping bounds across different $j$. For instance, in the left panel, students whose local first-best is Medicine at UChile are more likely to graduate from that program after assignment to the Bachillerato than those whose first-best is Medicine at Universidad de Santiago de Chile or Math at PUC. In the right panel, students with Medicine at UChile as their first-best are more likely to graduate from a top university when assigned to the Bachillerato than those with Math at PUC as their first-best.
The observed heterogeneity may arise from two sources. First, differences in preferences—potentially correlated with effort and ability—can influence graduation outcomes. Second, variation in academic preparation or ability, proxied by test scores, could also play a role as we are comparing subpopulations with $S_{j}=c_{j}$ for different $j$’s in Figure 3. We show that this second channel is unlikely to explain the differences observed in Figure 3. If admissions across programs $j$ were based on a single priority score, it would suffice to show that the cutoffs $c_{j}$ are close or that applicant score distributions are similar around them. However, priority scores vary by program, as each $S_{j}$ is a weighted sum of six primary components with program-specific weights. [^15] Primary scores: NEM, math, Spanish, science, history, and the max of science and history. Weights can be zero but always sum to 100%. Therefore, to assess comparability across cutoffs, we analyze the Euclidean distance between students’ six-dimensional primary score vectors. We show that at any of the cutoffs $c_{j}$ of interest, students within the bandwidth of the cutoff are, in terms of the Euclidean distance between their vectors of primary scores, on average as close to each other as they are to students around the other cutoffs of interest. Results are shown in Table C-3 and discussed in detail in Appendix C.
We therefore interpret the results in Figure 3 as evidence that students’ preferences matter for their graduation outcomes, suggesting that preferences are correlated with effort choices or ability that is not captured by test scores. [^16] Evidence from Heckman and Kautz (2012) shows that achievement tests do not fully capture soft skills or personality traits that influence labor market outcomes (see also Borghans et al. (2008), Almlund et al. (2011)). Structural models support this, showing that factors beyond test scores affect educational choices. For example, Arcidiacono (2005) find that, even conditional on SAT scores, some individuals have higher college admission chances, financial aid likelihood, labor market earnings without college, and returns to all majors.
Role of behavioral assumptions.
@@
We next examine how different behavioral assumptions affect the estimated bounds. Table C-4 in Appendix provides descriptive statistics and the estimated $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ for each pair $(j,k)$ and each set of assumptions. The different assumptions are used to construct $\bm{\mathrm{Q}}_{j}$ for each individual and therefore determine the conditioning set $\{\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ and the values of $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$. Under WPO, $\bm{\mathrm{Q}}_{j}$ is larger than under SPO, resulting in a larger conditioning set (as shown in Table C-4, Columns (1) vs. (4)). Given a fixed bandwidth around the cutoff $c_{j}$, the set $\{\omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\{(j,k)\}\neq\emptyset\}$ being larger under WPO than under SPO means that the denominator in the equations defining $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ is larger under WPO than under SPO. In addition, as the set $\bm{\mathrm{Q}}_{j}$ is larger when constructed under WPO than when constructed under SPO, the share of individuals for whom $Q_{j}$ is found to be a singleton is lower under WPO than it is under SPO. This means that the numerator in the equations defining $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ is smaller under WPO than under SPO. Overall, this leads to $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ being smaller under WPO than under SPO, as seen in Table C-4. The increase in $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ as we impose SPO instead of WPO unambiguously tends to reduce the width of the bounds. The exact extent to which the width of the bounds shrinks as we impose SPO instead of WPO also depends on the joint distribution of the outcome $Y$ and $\bm{\mathrm{Q}}_{j}$, however. Imposing UMAS eliminates certain elements from individuals’ $\bm{\mathrm{Q}}_{j}$s. As a consequence, following the same logic as above, imposing UMAS shrinks the conditioning set $\{\omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\{(j,k)\}\neq\emptyset\}$ and increases $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ (see Columns (3) vs. (4)). Again, the increased values of $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ are the key force that tends to reduce the width of the bounds as UMAS is imposed. The fields assumption, when added, further trims the set $\bm{\mathrm{Q}}_{j}$. Students that are part of $\{\omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\{(j,k)\}\neq\emptyset\}$ without the fields assumption but do not report programs from $\mathcal{F}(j)$ or $\mathcal{F}(k)$ are not part of $\{\omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\{(j,k)\}\neq\emptyset\}$ with the fields assumption. As before, this increases $\underline{\delta}_{j,k}^{+}$ and $\underline{\delta}_{j,k}^{-}$ (see Columns (2) vs. (3)).
Note, however, that the impact of the fields assumption depends on the overlap between fields of $j$ and $k$. Under stability, the first coordinate of $P_{j}$ and the first coordinate of any element of $\bm{\mathrm{Q}}_{j}$ coincide for students just above $c_{j}$, while the second coordinates of $P_{j}$ and any element of $\bm{\mathrm{Q}}_{j}$ coincide for students just below $c_{j}$ (Proposition 2). This means that, absent the fields assumption, each student in the set $\{\omega:\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ includes a program from either $\mathcal{F}(j)$ or $\mathcal{F}(k)$ in her reported preference. As a consequence, the fields assumption can only reduce the set $\{\omega:\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ for pairs $(j,k)$ for which $\mathcal{F}(j)\neq\mathcal{F}(k)$. In our empirical context, the Bachillerato de Ingreso Comun belongs to the STEM field. Therefore, the fields does not affect results for the pair with Math at PUC as $j$, since both $j$ and $k$ fall within STEM. However, for pairs where $j$ belongs to Health/Medicine, the assumption does help reduce the set $\{\omega:\bm{\mathrm{Q}}_{j}\cap\{(j,k)\}\neq\emptyset\}$ and narrow the bounds.
5.2.2 Treatment Effects
We now illustrate our method for estimating treatment effects, as opposed to average structural functions. Specifically, we estimate the effect on graduation outcomes of being assigned to Medicine at PUC Santiago instead of being assigned to a second-best option. We focus on this program for two main reasons. First, its high selectivity and popularity ensure a sufficiently large sample for precise estimation. Second, its competitiveness and highly predictable cutoff make it likely that applicants adjust their submissions strategically around the cutoff—a type of behavior our approach is designed to guard against. We consider the five most common next-best options to Medicine at PUC: Medicine at UChile, at Universidad de Concepción, and at Universidad de Santiago de Chile; and Science, and Engineering at PUC. As in earlier analyses, we construct our estimation sample using a 30-point bandwidth around the cutoff $c_{j}$.
Figure 4 presents the estimated bounds. The left panel shows treatment effects on the probability of graduating from Medicine at PUC Santiago; the right panel shows treatment effects on the probability of graduating from a top university. Starting with the left panel, we find, reassuringly, that assignment to Medicine at PUC increases the probability of graduating from that program. In all five cases, the bounds derived under WPO+UMAS+fields exclude zero, identifying a positive treatment effect. Importantly, the magnitude of the effect varies across next-best alternatives. For example, students whose second-best is Science at PUC appear to benefit less from admission to Medicine at PUC than those whose next-best is Medicine at UChile or at Universidad de Santiago de Chile. These differences are reflected in non-overlapping bounds across several cases.
The right panel reveals similar patterns. Admission to Medicine at PUC increases the probability of graduating from a top university in four of the five cases, and regardless of whether the next-best option is at a top university itself. For instance, the estimated bounds indicate a positive effect both for students whose next-best program is Science or Engineering at PUC and for those whose next-best is Medicine at Universidad de Concepción or Universidad de Santiago de Chile. There again, the bounds under the WPO+UMAS+fields assumption set highlight substantial heterogeneity across next-best alternatives.
As we relax assumptions, the bounds widen and become less informative—their ability to recover the sign of the treatment effect diminishes. This exercise thus helps clarify the identifying power of each assumption, providing researchers with a framework to evaluate the empirical content of different preference assumptions in similar settings.
Figure 4: Effects of assignment to Medicine at PUC vs. next-best alternatives
The figure shows estimated bounds for the average effect on the probability of graduation (y-axis) of assigning students to Medicine at PUC as opposed to assigning them to a second-best program. Bounds are shown for five different second-best programs (x-axis), two binary outcomes (left and right panels), and four behavioral assumptions used to construct $\bm{\mathrm{Q}}_{j}$ (four bars for each second-best program). The outcomes are graduation from Medicine at PUC on the left panel and graduation from a top university (PUC or UChile) on the right panel. The four behavioral assumptions are, from left to right, SPO, WPO in combination with UMAS and fields, WPO in combination with UMAS, and WPO alone. All bounds are estimated with a 30-point bandwidth on either side of the assignment cutoff of the program of interest (Medicine at PUC). Note that these are estimated bounds and not inference bounds.
5.2.3 On the Importance of Accounting for Strategic Behavior
Finally, to illustrate more broadly the relevance of our bounding approach, we show that naively ignoring strategic behavior in the identification of treatment effects can produce misleading results. In what follows, we compare our bounding approach to “naive estimates”, that is, estimates of
$\mu_{j,p}:=\mathbb{E}[Y\mid P_{j}=(j,k),S_{j}=c_{j}^{+}]-\mathbb{E}[Y\mid P_{j}=(j,k),S_{j}=c_{j}^{-}].$
We look at 302 pairs $(j,k)$ which we select as follows. First, we consider all pairs $(j,k)$ such that program $j$ is available on the platform at any point between 2004 and 2010 and at least one student lists $k$ right after $j$ in his ROL. Second, we count the number of students with score $S_{j}$ within a 30-point bandwidth of the year-specific $c_{j}$ who $P_{j}=(j,k)$. We find 327 pairs $(j,k)$ with a least 50 such students These pairs involve 137 distinct programs $j$, each of them associated with a number of $k$’s ranging from one to 18. We exclude the 25 pairs $(j,k)$ associated with three of these programs $j$ because no graduates are recorded in the data for these programs. This yields a final set of 302 pairs involving 134 distinct $j$’s.
As a simple procedure to construct a sample local to each admission cutoff $c_{j}$, we first use a 30-point bandwidth on either side of that cutoff. For a number of our $j$’s of interest, there exists another program $\ell$ that uses the same placement score as $j$ and whose cutoff $c_{\ell}$ is within a 30-point distance of $c_{j}$. In these cases, we use a bandwidth smaller than 30 around the cutoff $c_{j}$ of interest, so it does not include any such $c_{\ell}$, and the bandwidth is possibly different across sides of $c_{j}$.
Table 1 provides statistics on the estimated bounds we obtain for each of the two outcomes of interest. From the table, we can see that naively ignoring strategic behavior in the identification of treatment effects can produce misleading results. First, recall that the naive estimand $\mu_{j,p}$ identifies our causal parameter of interest if we assume everyone near $c_{j}$ is a truth-teller. One might expect naive estimands to always lie within our bounds since truth-telling is a special case of our behavioral assumptions. However, Table 1 shows this is not true. Under the SPO and assumptions, the naive estimates fall outside our bounds for 12% of the pairs of interest for our first outcome and 15% of the pairs of interest for our second outcome. The fact that naive estimates fall outside our bounds is consistent with strategic behavior of agents. Indeed, Part III of our SD example (Section 4.1, right after Proposition 2), shows that the distribution of the potential outcomes conditional on reported local preferences and the priority score can change sharply at the cutoff in the presence of strategic behavior. In such a case, the naive estimand $\mu_{j,p}$ no longer equals the average treatment effect and may fall outside our bounds. More broadly, this also suggests that caution is warranted in cases like the Chilean case where most but not all students submit $P$ with $|P|<K$ or very few students are assigned to their last-listed option. It may not be safe to simply assume that everyone is a truth-teller just because the cap does not bind for most people. In other words, naive estimates may be severely biased even if the share of constrained students is small.